OpenStudy (anonymous):

I need help bad, can someone please help me? a +7b = 13 a+ 2b = -2 there is on solution it is have to type a order pair. i got a+7b =13 a+ 2b =-2 a+7b =13 2+2= 4 b= 12? simply a = 3 ? can anyone help me?

OpenStudy (anonymous):

I will really appreciate if someone can help me?

OpenStudy (goten77):

hmm this is what ied do a+2b=-2 this means a = -2-b plug this into second and get -2-2b+7b=13 clean up and get 5b=15 then b=3 can u do the rest?

OpenStudy (anonymous):

okay 2b+7b=13 and 5b=15=3 so that mean that the pair are (5b=15, 3) right?

jimthompson5910 (jim_thompson5910):

Goten77 showed that b = 3 a +7b = 13 a +7(3) = 13 ... plug in b = 3 a + 21 = 13 a = 13 - 21 a = -8 -------------------------------- or a+ 2b = -2 a+ 2(3) = -2 a + 6 = -2 a = -2 - 6 a = -8 So either way, we get a = -8

jimthompson5910 (jim_thompson5910):

so the solution as an ordered pair is (-8, 3)

OpenStudy (anonymous):

why is it u do it so easy and the other person confused me? okay thank you so much, it like 7x 3= 21 then subtract 21 by 13 = -8 thanks

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (goten77):

well i thought u would follow the work....